White light with wavelength range 450 - 900 nm is incident normally on a thin film

of refractive index 1.3. Light of wavelengths 500, 600, 750 nm is reflected back.

What is the minimum thickness of the film? If we could use light from 400 - 4000

nm, what other wavelengths will be reflected back?


1
Expert's answer
2021-11-24T19:14:26-0500

"\Delta x=2\mu t=(2n-1)\frac{\lambda}2,"

"\lambda=\frac{4\mu t}{2n-1,}"

"t=\frac{(2n-1)\lambda}{4\mu}=96~nm,"

"\lambda':{410,~425,~460}."